Demand = 20-P; Supply = P/3; both linear. What is deadweight loss associated with a $4/unit output tax levied on consumers?


1
Expert's answer
2021-10-31T18:28:43-0400

Given

"Qd=20-p\\Qs=\frac{P}{3}"

Now we put unit tax of 4




Of=ps+4

With tax

"20-pd=\frac{ps}{3}\\20-ps-4=\frac{ps}{3}\\ps=\frac{16×3}{4}=122\\ps=16"

"Q=4"


Dead weight loss"=\frac{1}{2}×(5-4)(16-12) \\"

"=2"